Session: Nov 2010 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (Uwe Bach at Fri Nov 05 11:47:01 2010) Minor failures: ANT
Still some oil leakage at the elevation gear. Another stop to fix that. started later at 11:20 UT.

Medicina (Franco Fiocchi at Fri Nov 05 14:40:39 2010) Success:

Torun (K. Borkowski at Fri Nov 05 15:55:58 2010) Success:
No known problems.

Westerbork (Geert Kuper at Fri Nov 05 19:38:41 2010) Success:
Apart from RFI no known problems.

Onsala (Michael Lindqvist at Sat Nov 06 08:35:13 2010) Success:
RFI and rain. Otherwise no known problems.